Compute the zeroes of the polynomial x²-4x-8.Also establish a relationship between the zeroes and coefficient

Grade:10

1 Answers

Harshit Singh
2 years ago
Dear Student

x2-4x-8=0
by formula method
x= 4 + √48)/2 , (4-√48)/2
x= 2+2√3 , 2-2√3

sum of roots = -b/a = 4
product of roots = c/a = -8

Thanks